# Curvatures and Non-metricities in the Non-Relativistic Limit of Bosonic Supergravity

### v1

@@ Abstract
We construct a diffeomorphism-covariant formulation of the non-relativistic (NR) limit of bosonic supergravity. This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to the NR degrees of freedom. The construction is purely geometrical and is based on a torsionless connection. The non-metricities are associated with the gravitational fields of the theory, $\tau_{\mu\nu},h_{\mu\nu},\tau^{\mu\nu}$ and $h^{\mu\nu}$, and are fixed by requiring compatibility with the relativistic metric. We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature. Our results establish an equivalence between the proposed construction and the intrinsic torsion framework of string Newton–Cartan geometry. We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
@@ I Introduction
$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (8)
@@
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
@@
Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity. We expect this framework to serve as a useful starting point for higher-derivative extensions and for exploring duality-covariant formulations of string theory in the NR limit.
@@ III Non-metricities
$Q^{h^{-1}}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
@@ IV.1 The Riemann tensor
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{(4)}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{(2)}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
$+c^{(-2)}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{(-4)}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
@@
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\nu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
@@
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
@@
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\nu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
@@
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
@@
$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
@@ IV.2 The Ricci tensor
$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
@@
$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
@@
$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
@@ V.1 NR-Bosonic supergravity Lagrangian: covariant (re-)formulation
$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
$+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}$
$+2\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\nu]}\tau_{\mu\alpha}+\nabla_{\mu}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi$
$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
@@ V.2 Non-relativistic bosonic $\alpha^{\prime}$-corrections
$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}$ (53)
@@
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}$ (56)

### v2

@@ Abstract
We construct a diffeomorphism-covariant formulation of the non-relativistic (NR) limit of bosonic supergravity. This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to the NR degrees of freedom. The construction is purely geometrical and is based on a torsionless connection. The non-metricities are associated with the gravitational fields of the theory ($\tau_{\mu\nu}$, $h_{\mu\nu}$, $\tau^{\mu\nu}$, $h^{\mu\nu}$) and are fixed by requiring compatibility with the relativistic metric. We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature. Our results establish an equivalence between the intrinsic torsion framework of string Newton–Cartan geometry and the proposed construction and. We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
@@ I Introduction
$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta^{a^{\prime}b^{\prime}}\,$, (8)
@@
$\tau_{\mu}{}^{a}\tau^{\mu}{}_{b}$ $=\delta^{a}_{b}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
@@
Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity. We expect this framework to serve as a useful starting point for higher-derivative extensions in both supergravity and duality-covariant frameworks, in the NR limit of string theory.
@@ III Non-metricities
$Q^{(h^{-1})}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
@@ IV.1 The Riemann tensor
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{4}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{2}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
$+c^{-2}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{-4}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
@@
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\mu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
@@
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
@@
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\mu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
@@
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
@@
$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
@@ IV.2 The Ricci tensor
$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
@@
$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}\nabla_{\alpha}\tau_{[\mu\sigma}(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}\nabla_{\alpha}h_{[\mu\sigma}(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
@@
$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
@@ V.1 NR-Bosonic supergravity Lagrangian: covariant (re-)formulation
$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+h^{\mu\sigma}\tau^{\alpha\beta}\nabla_{\alpha}\tau_{[\mu\sigma}(\nabla_{\nu]}h_{\epsilon\beta}-\frac{1}{2}\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
$+\tau^{\mu\sigma}h^{\alpha\beta}\nabla_{\alpha}h_{[\mu\sigma}(\nabla_{\epsilon}\tau_{\nu]\beta}-\frac{1}{2}\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}+4\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\mu}\tau_{\nu]\alpha}-\frac{1}{2}\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}$
$+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$
$-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
@@ V.2 Non-relativistic bosonic $\alpha^{\prime}$-corrections
$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}\,.$ (53)
@@
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}\,.$ (56)